OpenStudy (anonymous):

Please help<3 Solve and graph the absolute value inequality: |4x + 1| ≤ 5. A number line with closed dots on −1.5 and 1 with shading going in the opposite directions. B number line with open dots on −1.5 and 1 with shading in between. C number line with closed dots on −1 and 1 with shading in between. D number line with closed dots on −1.5 and 1 with shading in between.

OpenStudy (phi):

if you have < do you use open or closed dots? if you have \( \le\) do you use open or closed dots?

OpenStudy (anonymous):

Thank you for your time. < closed? ≤ open? I think the answer is D? @phi

OpenStudy (phi):

yes. that helps rule out one of the choices. when you see absolute value | |, you should think "distance from a middle point" in other words, it will be an interval. That rules out A it is down to C or D. to choose, we have to figure out the end points. You can guess or figure it out.

OpenStudy (phi):

to figure it out we break the problem into 2 parts: if what is inside the | | is positive, we can drop the | | so let's assume the inside is positive: \( |4x + 1| ≤ 5 \) becomes \( 4x + 1 ≤ 5 \) and if we solve we get \( x \le 1\) if the stuff inside is negative, the absolute value makes it positive. We can have the same effect by multiplying by -1 and dropping the | | \( -(4x + 1) ≤ 5 \) to solve we have to be careful. if we multiply or divide by a negative, it changes the order of the <. example -2 < -1 but -1*-2 < -1*-1 gives 2 < 1, which is wrong. one way around this is always add or subtract. example: add 4x+1 to both both sides \[ -(4x+1)+(4x+1) ≤ 5 + (4x+1) \\ 0 \le 4x+6 \\ -6 \le 4x \\ -1.5 \le x \] that gives the lower point.

OpenStudy (anonymous):

-6<=4x<=4 -1.5<=x<=1